The pistons in an automobile engine move up and down in the cylinders. If t is in milliseconds, the distance from the top of the piston to the top of the cylinder is given in centimeters by D=7+6 sin (4m)




(a) identify the amplitude, period and frequency of function




(b) Find the largest and the smallest clearance between the piston and the top of the cylinder.

1
Expert's answer
2022-02-07T16:14:00-0500

(a) Amplitude "A=6."

Period "P=\frac{2\pi}{4}=\frac{\pi}{2}."

Frequency "f=\frac{2\pi}{P}=4."


(b) Largest distance: "7+6=13" cm.

Smallest distance: "7-6=1" cm.